OpenStudy (anonymous):

WILL FAN AND MEDAL What is the ratio of x to y? 3/2=2x/2x+y

OpenStudy (michele_laino):

hint: \[\frac{{2x + y}}{{2x}} = \frac{2}{3}\]

OpenStudy (michele_laino):

no, please another hint: \[\frac{{2x + y}}{{2x}} = 1 + \frac{y}{{2x}} = 1 + \frac{1}{2}\left( {\frac{y}{x}} \right)\]

OpenStudy (michele_laino):

find the ratio y/x first

OpenStudy (anonymous):

oh ok thx

OpenStudy (michele_laino):

thanks!